OpenStudy (anonymous):

find the measure of an angle if its supplement measures 12 less than three times its complement

OpenStudy (anonymous):



OpenStudy (anonymous):

supplement of angle 180 -x 12 less than three times its complement: 3(90 -x) -12 => 180 -x = 3(90 -x) -12 3x - x = 270 - 180 -12 Can you continue?

OpenStudy (anonymous):

to be honest i got it till the point were u stoped cause ur explaining it to me what should u do next

OpenStudy (anonymous):

3x -x = ...?

OpenStudy (anonymous):

2x

OpenStudy (anonymous):

270 -180 -12 = ...?

OpenStudy (anonymous):

178

OpenStudy (anonymous):

were did the 90come from'

OpenStudy (anonymous):

angle complement: 90 -x

OpenStudy (anonymous):

so i can divide the 2 into 178

OpenStudy (anonymous):

39

OpenStudy (anonymous):

huh

OpenStudy (anonymous):

78/2=39

OpenStudy (anonymous):

like about this problem

OpenStudy (anonymous):

no i understood it better thanks